XYZ has the following financial information for 2009:

Sales =$ 2 M, Net Inc. =$ 0.4 M, Div. =$ 0.1 M

C.A. =$ 0.4 M, F. A .=$ 3.6 M

C.L. =$ 0.2 M, LTD=$ 1 M, C.S .=$ 2 M, R.E. = $ 0.8 M


What is the sustainable growth rate?

If 2010 sales are projected to be $ 2.4 M, what is the amount of external financing needed, assuming X Y Z is operating at full capacity, and profit margin and payout ratio remain constant?
Answer #1

a). ROE = net income / shareholders’ equity = $.4M / ($2M + $.8M) = .1429

Payout ratio = dividends/net income = .1M/.4M = .25

Plowback ratio (b) 1 – payout ratio = 1 - .25 = .75

Sustainable growth rate = ROE X b / 1 – ROE X b = .1429 X .75 / (1 – (.1429 X .75)) = .12


b). Profit margin = net income/sales = .4M/2M = .2

Projected net income = profit margin X projected sales = .2 X $2.4M = $.48M

Projected addition to retained earnings = projected net income X (1 – payout ratio)

= $.48M X (1-.25)= $.48M X .75 = $.36M%

change in sales = ($2.4M - $2M)/$2M = .22009

total assets = $.4M + $3.6M = $4M

Projected total assets = $4M X 1.2 = $4.8M

Projected C.L. = $.2M X 1.2 = $.24M

Projected R.E. = 2006 R.E. + projected addition to R.E. = $.8M + $.36M = $1.16M

Projected liabilities and owners’ equity = projected C.L. + LTD + C.S. + projected R.E.

= $.24M + $1M + $2M + $1.16M = $4.4M

EFN = Increase in Assets - Increase in Liabilities - Increase in Retained Earnings

= $0.8M - $0.04M - $0.36M = $0.4M


answered by: Mon